# Joint AP-UE Association and Power Factor Optimization for Distributed Massive MIMO

### v3

@@ I Introduction
A significant amount of work has been undertaken to address power control and AP-UE association problems in distributed mMIMO systems. However, it is worth noting that much of it tends to focus entirely on either power control [3, 5, 8, 6], or AP-UE association [6, 9, 10]. In [5] and [11], the authors focus on a downlink system, where the inherent coupling between AP-UE association and power control makes the joint analysis less complex in contrast to the uplink scenario, which lacks such natural coupling between the two. This highlights the importance of customized methodologies for extending joint optimization techniques from downlink to uplink distributed mMIMO systems. While [12] introduces a deep learning-based joint approach for power control and AP-UE association in uplink scenarios. It has high computational cost of extensive training and lacks the explainability associated with traditional models. In [13] and [14], a novel approach is adopted by formulating an optimization problem from the perspective of maximizing the minimum Signal-to-Interference-Noise ratio for joint power control and AP-UE association weights for the uplink distributed mMIMO. However, there is an inherent difficulty in allocating APs to UEs within the defined optimization challenges and it prompted the authors to suggest different methods for AP-UE scheduling.
@@ II System Model
We consider a system consisting of $T$ single antenna UEs and $M$ multi-antenna APs, where each AP has $A$ antennas such that $T<<MA$. The APs and UEs are distributed uniformly over the given coverage area and APs are jointly serving all users over the same time and frequency resources. In order to help in the signal processing and coordination among APs, they are connected with a CPU via a front-haul connection. We consider the TDD mode of operation where the AP-UE channel vector is estimated using the uplink transmission. To model our channel, we have considered the block fading model with a coherence block of length $L_{c}$. Let $L_{p}$ be the length of orthogonal pilots such that $L_{p}$ length of coherence block is used for uplink pilot training. The Rayleigh fading coefficient representing the small-scale fading is denoted by $\textbf{h}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, whereas the large scale-fading coefficient (LSFC) that incorporates both path-loss and shadowing effects is denoted by $\beta_{mt}$. The channel vector $\textbf{g}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, that incorporates both the small-scale fading as well as large scale fading, is defined as $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$. We assume that $\textbf{h}_{t,m}{\sim}\mathcal{N}_{\mathbb{C}}(0,1)$ is independent and identically distributed (i.i.d.) random vector.
@@ II-A Uplink Pilot Training
We assume that during uplink channel estimation, each UE $t\in\mathcal{T}\overset{\Delta}{=}\{1,...,T\}$ transmits the pilot sequence $\sqrt{p_{p}L_{p}\eta_{t}}\mathbf{\Psi}_{t}\in\mathbb{C}^{L_{p}\times 1}$, such that $||\mathbf{\Psi}_{t}||^{2}=1$, where $p_{p}$ represents the normalized signal-to-noise power. The received signal $\textbf{y}^{pilot}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ at the $m^{th}$ AP is given by :
$\textbf{y}^{pilot}_{m}=\sum_{t\in\mathcal{T}}\sqrt{p_{p}L_{p}\eta_{t}}\textbf{g}_{mt}\mathbf{\Psi}_{t}+\textbf{n}_{m},$
@@
$\hat{\textbf{g}}_{mt}=\frac{\sqrt{p_{p}L_{p}\eta_{t}}\beta_{mk}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\eta_{t}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t}.$
@@
$\gamma_{mt}=\left(\frac{p_{p}L_{p}\eta_{t}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\eta_{t}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\right).$
@@ II-B Uplink Data Transmission
$\hat{\textbf{y}}^{u}_{t}=\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{y}^{u}_{m}]_{a},$
where [.$[.]_{a}$ represents the $a^{th}$ element of the vector [.][.][ . ]. Decomposing the $\hat{\textbf{y}}^{u}_{t}$ as in [3], results in:
$\hat{\textbf{y}}^{u}_{t}=\text{DS}_{t}\ x_{t}+\text{BU}_{t}\ x_{t}+\sum_{t^{\prime}=1,t^{\prime}\neq t}^{T}\text{UI}_{tt^{\prime}}\ x_{t^{\prime}}+n_{t},$
@@
Treating interference as effective-noise, we get SE for the UE $t$ as:
@@
$\Gamma^{*}_{t}\overset{\Delta}{=}\frac{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{{I}_{t}},$ (2)
@@
$\begin{split}{I}_{t}\overset{\Delta}{=}\sum\limits_{t^{\prime}=1,t^{\prime}\neq t}^{T}\!\!\!\!\!A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}(\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}})^{2}+\sum\limits_{t^{\prime}=1}^{T}\!\!\!Ap_{u}\eta_{t^{\prime}}^{u}\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!Ad_{mt}\gamma_{mt},\end{split}$ (3)
@@ III Proposed Joint Solution using Relaxation Technique
$\begin{split}&\underset{\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-\frac{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right),\\ \end{split}$ (9a)
@@
$\begin{split}\underset{\bm{\eta^{u}},\textbf{D},\mathbf{\Gamma},\textbf{u}}{\max~{}}&\ \ \sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-2u_{t}\sqrt{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}\right)\\ &+\sum\limits_{t=1}^{T}u_{t}^{2}\left({A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right),\end{split}$ (10)
where $\textbf{u}=[u_{1},u_{2},...u_{T}]^{\intercal}$.
@@
After substituting the optimal value of u in equation (10) and then solving optimization problem (12) is equivalent to solving Problem (9). Notice that Problem (12) is a convex optimization problem in different blocks. If we fix D, $\mathbf{\Gamma}$ and u, then the objective function of problem (12a) becomes concave with respect to $\bm{\eta^{u}}$. Then, the problem reduces to:
@@
Similarly, upon fixing $\bm{\eta^{u}}$, $\mathbf{\Gamma}$ and u, the objective function of (12) becomes concave with respect to D. Then, the optimization problem becomes:
@@
Problem 12 can be solved following the proposed iterative approach as follows. This problem is decomposed into two distinct blocks, denoted as Problem 13 and Problem 14. Initially, the solution for Problem 13 is obtained while keeping D, $\mathbf{\Gamma}$, and u fixed. Subsequently, based on the solution derived for Problem 13, the solution for Problem 14 is determined. This is repeated until the convergence criterion is met. Algorithm 1 provides the details of this approach.
1:  Feasible Initialization: ${\bm{\eta_{t}^{u}}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall t$.
@@
7:     Solve optimization problem in (13) for ${\bm{\eta_{k}^{u}}}^{(i)}$.
@@
11:     Solve optimization problem in (14) for $\textbf{D}^{(i)}$.
@@
For the convergence proof of the proposed Algorithm 1, please refer to the appendixV.
@@ IV Performance Evaluation
We performed numerical simulations on a $1000\times 1000$ square meters geographical area with APs and user equipment UEs randomly distributed. To simulate an infinite area network, a wrapped-around approach is used [6]. Large-scale fading coefficients are modeled using a three-slope approach for path loss, incorporating an 8 dB standard deviation for shadow fading. The minimum QoS parameter is taken as 0.1 bits/s/Hz for each UE. Throughout our simulations, we keep $T=30$, $M=50\ \text{or}\ 100$, $A=4$, $\alpha=0.0005\ \text{or}\ \alpha=0.001$, $L_{c}=200$ and $L_{p}=5$, while other parameters stay exactly the same as those outlined in [3] or as explicitly stated. All these numerical results are averaged over $500$ simulations.
@@
Figure 1: The variation of maximum fronthaul load on an AP, objective function, and the sum SE, when $M=50$.
@@
Figure 1 shows the variations in maximum front-haul load, objective function, and the sum SE for two different regularization coefficient values, with the number of antennas fixed at $50$. This highlights how the penalty function affects the network’s maximum front-haul load and total sum-throughput. Notably, the objective function value consistently remains lower than the total SE value, exhibiting a drop attributed to the penalty function. When the regularization coefficient is increased by two times, the maximum front-haul load on an AP decreases by $14\%$. This reduction is due to the increased regularization coefficient, which limits AP-UE associations. As a consequence of the lesser number of AP-UE associations, there is a reduction in the sum SE by nearly $1\%$. Similarly, Figure 2 illustrates variability in maximum front-haul load, objective function, and sum SE for two different regularization coefficient values, however with $M$ set to 100. For this scenario too, when the regularization coefficient is increased two times, the maximum front-haul load and total SE show a similar decreasing trend as in Fig. 1. Therefore, the proposed approach effectively decreases the maximum front-haul load while not significantly reducing overall sum throughput, and this is achieved by a careful tuning of regularization coefficient.
@@
Figure 2: The variation of maximum front-haul load on an AP, objective function, and the sum SE, when $M=100$.
@@
Examining Figures 1 and 2, we observe a noteworthy trend as the number of antennas ($M$) increases from 50 to 100. The maximum front-haul load increases by $14\%$ and $17\%$ with regularization coefficients ($\alpha$) of 0.0005 and 0.001, respectively. This growth can be attributed to the growing number of APs in the fixed area. As the AP count increases, there are more APs in the vicinity of each UE, resulting in a higher percentage of APs offering a strong channel or a larger large-scale fading coefficient for each UE. This gainful influence of numerous APs on the large-scale fading coefficient causes a change in the influence of the penalty function, allowing more APs to associate with each UE. While this increased association raises the maximum front-haul load, it also results in a significant ($30\%$) boost in the overall sum-throughput. In essence, the system responds to the increased AP count by increasing the strength of channels across numerous APs, hence increasing the network’s overall throughput. The penalty function is a dynamic component of this process, optimizing AP-UE associations to optimize the objective function while reducing negative affects via penalties.
@@
Figures 3 and 4 illustrate the variation of sum SE and per-UE SE respectively, of the proposed solution in comparison to the following four scenarios: a) where UEs transmit at full power and are served by all APs; b) fractional power control [8]; c) where only power control is performed using the proposed algorithm; and d) where only AP-UE association is performed using the proposed algorithm.
Figure 3 shows that when AP-UE association and power control are carried out simultaneously, the achievable sum SE is the largest. The proposed solution’s sum SE surpasses that of scenarios ’a’, ’b’,’c’ and ’d’ by $16.5\%$, $4.3\%$, $1.5\%$ and $2\%$, respectively, when $M=50$. Additionally, for case with $M=100$, the proposed solution outperforms these scenarios by $11\%$, $3\%$, $1.3\%$, and $1.6\%$, respectively. This demonstrates the effectiveness of the proposed technique in optimizing the sum SE by jointly addressing the AP-UE association and power control problems. Notably, when only AP-UE association or only power control is conducted with the proposed method, the sum SE sees a improvement when compared to the other two scenarios ’a’ and ’b’. This improvement occurs while preserving the minimum QoS limitation. The proposed approach’s versatility and effectiveness are demonstrated by its ability to improve network performance even when improving just a single aspect of the system, like AP-UE association or power control.
Indeed, our proposed technique aims to enhance the sum SE. Figure 4 shows how our strategy has a significant influence on the $90\%-likely$ per-user SE. Our proposed scheme outperforms scenarios ’a’, ’b’, ’c’, and ’d’ in terms of $90\%-likely$ per-user SE by $175\%$, $22\%$, $4.7\%$, and $11.7\%$, respectively, at $M=50$. Similarly, for $M=100$, our approach outperforms the same cases by $52.3\%$, $6.7\%$, $2.1\%$, and $3.2\%$ in terms of $90\%-likely$ per-user SE. The findings illustrate an additional benefit of the method we propose, demonstrating its massive contribution to improve per-user throughput while also increasing the total throughput.
@@
In Figure 5, we provide a numerical justification for relaxation proposed in (5). It shows the variation of the SE with and without the discretization of matrix D. The figure shows that for $m=50$, the variation in the SE is negligible with and without the discretization of elements of matrix D. For the case when $m=100$, the difference in SE is mere $0.2\%$. The negligible SE variation can be attributed to the fact that penalty parameter does not force elements of matrix D too much away from zero or one, thus our proposed solution does not require additional algorithm for AP-UE association.

### v4

@@ I Introduction
A significant amount of work has been undertaken to address power control and AP-UE association problems in distributed mMIMO systems. However, it is worth noting that much of it tends to focus entirely on either power control [3, 5, 8, 6], or AP-UE association [6, 9, 10]. In [5] and [11], the authors focus on a downlink system, where the inherent coupling between AP-UE association and power control makes the joint analysis less complex in contrast to the uplink scenario, which lacks such natural coupling between the two. This highlights the importance of customized methodologies for extending joint optimization techniques from downlink to uplink distributed mMIMO systems. While [12] introduces a deep learning-based joint approach for power control and AP-UE association in uplink scenarios. It has high computational cost of extensive training and lacks the explainability associated with traditional models. In [13] and [14], a novel approach is adopted by formulating an optimization problem from the perspective of maximizing the minimum Signal-to-Interference-Noise ratio for joint power control and AP-UE association weights for the uplink distributed mMIMO. However, there is an inherent difficulty in allocating APs to UEs within the defined optimization challenges. The solutions obtained provide the weights of the AP-UE association matrix and it prompted the authors to proposed additional algorithm for AP-UE scheduling.
@@ II System Model
We consider a system consisting of $T$ single antenna UEs and $M$ multi-antenna APs, where each AP has $A$ antennas such that $T<<MA$. The APs and UEs are distributed uniformly over the given coverage area and APs are jointly serving all users over the same time and frequency resources. In order to help in the signal processing and coordination among APs, they are connected with a CPU via a front-haul connection. We consider the TDD mode of operation where the AP-UE channel vector is estimated using the uplink transmission. To model our channel, we have considered the block fading model with a coherence block of length $L_{c}$. Let $L_{p}$ be the length of orthogonal pilots such that $L_{p}$ length of coherence block is used for uplink pilot training. The Rayleigh fading coefficient representing the small-scale fading is denoted by $\textbf{h}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, whereas the large scale-fading coefficient (LSFC) that incorporates both path-loss and shadowing effects is denoted by $\beta_{mt}$. The channel vector $\textbf{g}_{mt}\in\mathbb{C}^{\textit{A}\times 1}$, that incorporates both the small-scale fading as well as large scale fading, is defined as $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$. We assume that $\textbf{h}_{mt}{\sim}\mathcal{N}_{\mathbb{C}}(0,1)$ is independent and identically distributed (i.i.d.) random vector.
@@ II-A Uplink Pilot Training
We assume that during uplink channel estimation, each UE $t\in\mathcal{T}\overset{\Delta}{=}\{1,...,T\}$ transmits the pilot sequence $\sqrt{p_{p}L_{p}}\mathbf{\Psi}_{t}\in\mathbb{C}^{L_{p}\times 1}$, such that $||\mathbf{\Psi}_{t}||^{2}=1$, where $p_{p}$ represents the normalized signal-to-noise power. The received signal $\textbf{y}^{pilot}_{m}\in\mathbb{C}^{\textit{A}\times L_{p}}$ at the $m^{th}$ AP is given by :
$\textbf{y}^{pilot}_{m}=\sqrt{p_{p}L_{p}}\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\mathbf{\Psi}_{t}^{\mathit{H}}+\textbf{n}_{m},$
@@
$\hat{\textbf{g}}_{mt}=\frac{\sqrt{p_{p}L_{p}}\beta_{mk}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\textbf{y}^{pilot}_{m}\mathbf{\Psi}_{t}.$
@@
$\gamma_{mt}=\left(\frac{p_{p}L_{p}\beta_{mt}^{2}}{\sum_{t^{\prime}=1}^{T}p_{p}L_{p}\beta_{mt^{\prime}}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}+1}\right).$
@@ II-B Uplink Data Transmission
$\hat{{y}}^{u}_{t}=\sum_{m\in\mathcal{M}_{t}}\sum_{a=1}^{A}[\hat{\textbf{g}}_{mt}^{*}]_{a}[\textbf{y}^{u}_{m}]_{a},$
where [.$[.]_{a}$ represents the $a^{th}$ element of the vector [.][.][ . ]. Decomposing the $\hat{{y}}^{u}_{t}$ as in [3], results in:
$\hat{{y}}^{u}_{t}=\text{DS}_{t}\ x_{t}+\text{BU}_{t}\ x_{t}+\sum_{t^{\prime}=1,t^{\prime}\neq t}^{T}\text{UI}_{tt^{\prime}}\ x_{t^{\prime}}+n_{t},$ (1)
@@
By treating the sum of second, third and fourth terms in equation (1) as ”effective-noise” [3], we get uplink SE for the UE $t$ as [3]:
@@
$\Gamma^{*}_{t}\overset{\Delta}{=}\frac{A^{2}p_{u}\eta_{t}^{u}\left(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt}\right)^{2}}{{I}_{t}},$ (3)
@@
$\begin{split}{I}_{t}\overset{\Delta}{=}\sum\limits_{t^{\prime}=1,t^{\prime}\neq t}^{T}\!\!\!\!\!A^{2}p_{u}\eta_{t^{\prime}}^{u}\left|\Psi^{H}_{t}\Psi_{t^{\prime}}\right|^{2}\left(\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\frac{\beta_{mt^{\prime}}}{\beta_{mt}}\right)^{2}+\sum\limits_{t^{\prime}=1}^{T}\!\!\!Ap_{u}\eta_{t^{\prime}}^{u}\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!d_{mt}\gamma_{mt}\beta_{mt^{\prime}}+\!\!\!\sum\limits_{m\in\mathcal{M}}\!\!\!Ad_{mt}\gamma_{mt},\end{split}$ (4)
@@ III Proposed Joint Solution using Relaxation Technique
$\begin{split}&\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-\frac{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}{A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right),\\ \end{split}$ (10a)
@@
$\begin{split}\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}\ &\sum\limits_{t=1}^{T}\left(w\text{log}_{2}(1+\Gamma_{t})-\alpha||{\textbf{D}_{t}}||_{1}\right)+\sum\limits_{t=1}^{T}u_{t}^{2}\left({A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}+{I}_{t}}\right)\\ &-\sum\limits_{t=1}^{T}\left(\Gamma_{t}-2u_{t}\sqrt{w^{\prime}(1+{\Gamma}_{t})A^{2}p_{u}\eta_{t}^{u}(\sum\limits_{m\in\mathcal{M}}d_{mt}\gamma_{mt})^{2}}\right),\end{split}$ (11)
where $\textbf{u}=[u_{1},u_{2},...u_{T}]^{\intercal}$, $u_{t}$ is also the quadratic convex approximation parameter and need to be updated on the basis of solution of previous iteration.
@@
After substituting the optimal value of u in equation (11) and then solving optimization problem (13) is equivalent to solving Problem (10). Notice that Problem (13) is a convex optimization problem in different blocks. If we fix D, then the objective function of problem (13a) becomes concave with respect to $\bm{\eta^{u}}$. Then, the problem reduces to:
@@
Similarly, upon fixing $\bm{\eta^{u}}$, the objective function of (13) becomes concave with respect to D. Then, the optimization problem becomes:
@@
Problem 13 can be solved following the proposed iterative approach as follows. This problem is decomposed into two distinct blocks, denoted as Problem 14 and Problem 15. Initially, the solution for Problem 14 is obtained while keeping D fixed. Subsequently, based on the solution derived for Problem 14, the solution for Problem 15 is determined. This is repeated until the convergence criterion is met. Algorithm 1 provides the details of this approach.
1:  Feasible Initialization: ${\bm{\eta^{u}}}^{(0)}$, $\textbf{D}^{(0)}$ and $i=0$, $\forall t$.
@@
7:     Solve optimization problem in (14).
@@
11:     Solve optimization problem in (15).
@@
Convergence Analysis : For the convergence proof of the proposed Algorithm 1, please refer to the appendixV.
Complexity Analysis : Problem 14 has $T$ scalar variables, $3T$ linear constraints, so the complexity for solving Problem 14 is $\mathcal{O}(\sqrt{3T}4T^{3})$ [18]. Similarly, the complexity to solve Problem 15 with $MT$ scalar variables and $2(MT+T)$ linear constraint is give by $\mathcal{O}(\sqrt{2(MT+T)}(3MT+2T)(MT)^{2})$ [18].
@@ IV Performance Evaluation
We performed numerical simulations on a $1000\times 1000$ square meters geographical area with APs and user equipment UEs randomly distributed. To simulate an infinite area network, a wrapped-around approach is used [6]. Large-scale fading coefficients are modeled using a three-slope approach for path loss, incorporating an 8 dB standard deviation for shadow fading. The minimum QoS parameter is taken as $0.2$ bits/s/Hz for each UE. Throughout our simulations, we keep $T=40$, $M=100\ \text{or}\ 150$, $A=4$, $L_{c}=200$ and $L_{p}=5$, while other parameters stay exactly the same as those outlined in [3] or as explicitly stated. All these numerical results are averaged over $100$ simulations. We opted not to compare our proposed approach with works in [13] and [14]. As our proposed work focuses on the sum-throughput maximization whereas aforementioned works concentrate on maximizing the minimum Signal-to-Interference-Noise ratio. Furthermore, our proposed approach guarantees a minimum spectral efficiency of 0.1 bits/s/Hz for each UE, which was not explicitly addressed in above studies.
@@
Figure 1: The variation of maximum fronthaul load on an AP, objective function, and the sum SE, when $M=100$.
@@
Figure 1 shows the variations in maximum front-haul load, objective function, and the sum SE for two different regularization coefficient values, with the number of antennas fixed at $100$. This highlights how the penalty function affects the network’s maximum front-haul load and total sum-throughput. Notably, the objective function value consistently remains lower than the total SE value, exhibiting a drop attributed to the penalty function. When the regularization coefficient is increased by two and four times, the maximum front-haul load on an AP decreases by $13\%$ and $25\%$ respectively. This reduction is due to the increased regularization coefficient, which limits AP-UE associations. As a consequence of the lesser number of AP-UE associations, there is a reduction in the sum SE by nearly $0.3\%$ and $0.6\%$ respectively. Similarly, Figure 2 illustrates variability in maximum front-haul load, objective function, and sum SE for two different regularization coefficient values, however with $M$ set to $150$. For this scenario too, when the regularization coefficient is increased two and four times, the maximum front-haul load and total SE show a similar decreasing trend as in Fig. 1. Therefore, the proposed approach effectively decreases the maximum front-haul load while not significantly reducing overall sum throughput, and this is achieved by a careful tuning of regularization coefficient.
@@
Figure 2: The variation of maximum front-haul load on an AP, objective function, and the sum SE, when $M=150$.
@@
Examining Figures 1 and 2, we observe a noteworthy trend as the number of APs ($M$) increases from 100 to 150. The maximum front-haul load increases by $7.8\%$, $8.3\%$ and $8.4\%$ with regularization coefficients ($\alpha$) of 0.0005, 0.001 and 0.002, respectively. This growth can be attributed to the growing number of APs in the fixed area. As the AP count increases, there are more APs in the vicinity of each UE, resulting in a higher percentage of APs offering a strong channel or a larger large-scale fading coefficient for each UE. This gainful influence of numerous APs on the large-scale fading coefficient causes a change in the influence of the penalty function, allowing more APs to associate with each UE. While this increased association raises the maximum front-haul load, it also results in a significant ($15\%$) boost in the overall sum-throughput. In essence, the system responds to the increased AP count by increasing the strength of channels across numerous APs, hence increasing the network’s overall throughput. The penalty function is a dynamic component of this process, optimizing AP-UE associations to optimize the objective function while reducing negative affects via penalties.
@@
Figures 3 and 4 illustrate the variation of sum SE and 90%-likely SE respectively, of the proposed solution in comparison to the following four scenarios: a) where UEs transmit at full power and are served by all APs; b) fractional power control, when each AP serves all the UEs [8]; c) where only power control is performed using the proposed algorithm, when each AP serves all the UEs; and d) where only AP-UE association is performed using the proposed algorithm.
Figure 3 shows that when AP-UE association and power control are carried out simultaneously, the achievable sum SE is the largest. The proposed solution’s sum SE surpasses that of scenarios ’a’, ’b’,’c’ and ’d’ by $12\%$, $4\%$, $9\%$ and $5.5\%$, respectively, when $M=100$. Additionally, for case with $M=150$, the proposed solution outperforms these scenarios by $9\%$, $3.5\%$, $7\%$, and $4\%$, respectively.Our proposed approach outer-performs the scenarios ’a’ and ’b’ despite their failure to meet the minimum QoS requirement. Moreover, in the scenarios ’a’, ’b’ and ’c’, there is no regulation in front-haul load as all UEs are served by all APs. This demonstrates the effectiveness of the proposed technique in optimizing the sum SE by jointly addressing the AP-UE association and power control problems.
Indeed, our proposed technique aims to enhance the sum SE. Figure 4 shows how our strategy has a significant influence on the $90\%-likely$ per-user SE. Our proposed scheme outperforms scenarios ’a’, ’b’, ’c’, and ’d’ in terms of $90\%-likely$ per-user SE by $71\%$, $15\%$, $54\%$, and $45\%$, respectively, at $M=100$. Similarly, for $M=150$, our approach outperforms the same cases by $38\%$, $9\%$, $28\%$, and $18\%$ in terms of $90\%-likely$ per-user SE. The findings illustrate an additional benefit of the method we propose, demonstrating its massive contribution to improve per-user throughput while also increasing the total throughput.
@@
In Figure 5, we provide a numerical justification for relaxation proposed in (6). It shows the variation of the SE with and without the discretization of matrix D. The figure shows that for $M=100$ as well for $M=150$, the variations in the SE is negligible with and without the discretization of elements of matrix D. The negligible SE variations can be attributed to the fact that penalty parameter does not force elements of matrix D too much away from zero or one, thus our proposed solution does not require additional algorithm for AP-UE association.